# Efficient Policy Optimization in Robust Constrained MDPs with Iteration Complexity Guarantees

### v1

@@ 1 Introduction
[…]
@@ 2 Problem Formulation
$V^{\pi,P}_{c_{i}}(s):=\mathbb{E}_{P,\pi}\left[\underset{t=1}{\overset{\infty}{\Sigma}}\gamma^{t-1}\pi(a|s)c_{i}^{t}(s,a)\right],$ (4)
@@
[…]
@@
[…]
@@ 3.2 Policy Optimization Algorithm
$\pi_{t+1}\in\arg\min_{\pi\in\Pi}\langle\nabla_{\pi_{t}}J_{i}(\pi_{t}),\pi-\pi_{t}\rangle+\dfrac{1}{\alpha_{t}}\mathrm{KL}(\pi||\pi_{t})$ where $\text{where }i=\arg\max\{\frac{J_{c_{0}}^{\pi}}{\lambda},\{J_{c_{n}}^{\pi}-b_{n}+\xi\}_{n=1}^{K}\}$
where $\mathrm{KL}$ is the usual Kullback-Leibler divergence. Note that this is a convex optimization problem, and can be optimized efficiently. If we use, $\ell_{2}$ regularization, i.e., $||\pi-\pi_{t}||_{2}^{2}$, then it becomes a robust projected policy gradient (RPPG) adapted from the unconstrained version (see Appendix D[46, 36], a variant of which is used in [8] to find optimal policy for each $b_{0}$. Of course, our approach also works for $\ell_{2}$ norm which we define in Algorithm 4. Empirically, we observe that KL-divergence has a better performance, and provide iteration complexity for RNPG.
@@ 4 Theoretical Results
Thus, Algorithm 1 has an iteration complexity of $O(\epsilon^{-2})$. Also, the policy is feasible and only at most $\epsilon$-sub optimal. Our result improves upon the result $O(\epsilon^{-4})$ achieved in [8]. Further, they did not guarantee feasibility of the policy (rather, only $\epsilon$-violation). More importantly, we do not need to employ a binary search algorithm. Thus, our algorithm is computationally more efficient. Our dependence on $S$, $A$, and $(1-\gamma)^{-1}$ are significantly better compared to [8] as well. Note that for the unconstrained case the iteration complexity is $O(\epsilon^{-1})$ [45]; whether we can achieve such a result for robust CMDP has been left for the future.
@@ 5 Experimental Results
We evaluate our algorithms [^2] The complete implementation is available at https://github.com/VocenInquisitor/RNPG.git on two environments: (i) Garnet, and (ii) Constrained Riverswim (CRS). Additional experimental results are provided in Appendix E.
@@ 6 Discussions and Limitation
Note that since we are not assuming strict feasibility of the optimal policy, we can only bound the violation up to $\epsilon$. The key here is to use $\lambda=2H/\epsilon$ as we do not know $\xi$, and then obtain an $\epsilon^{2}$-close result using Lemma 4.2. This makes the iteration complexity of $O(\epsilon^{-4})$. Note that our dependence on $S$, $A$, and $1/(1-\gamma)$ are significantly better compared to EPIRC-PGS[8]. Further, we do not employ binary search. The proof is in Appendix C.
@@ Appendix A Proof of Proposition 1
$\langle Q_{s,i_{t}}^{\pi_{t},p_{t}},\pi_{t,s}-\pi_{t+1,s}\rangle-\dfrac{1}{\alpha}B(\pi_{t+1},\pi_{t})\leq||q_{s,i_{t}}^{\pi_{t},p_{t}}||_{\infty}||\pi_{t,s}-\pi_{t+1,s}||_{1}-\dfrac{1}{2\alpha}||\pi_{t,s}-\pi_{t+1,s}||_{1}^{2}$
$=\dfrac{-1}{2\alpha}(\alpha_{t}||Q_{s,i_{t}}^{\pi_{t},p_{t}}||_{\infty}-||\pi_{t,s}-\pi_{t+1,s}||_{1})^{2}+\dfrac{\alpha}{2}||Q_{s,i_{t}}^{\pi_{t},p_{t}}||_{\infty}^{2}$
$\leq\dfrac{\alpha}{2}||Q_{s,i_{t}}^{\pi_{t},p_{t}}||_{\infty}^{2}.$ (25)
@@
Here, we use the fact that $||Q_{s,a,c_{i}}^{\pi_{t},p_{t}}||_{\infty}\leq\dfrac{1}{1-\gamma}$. This is easy to discern for $i=\{1,\ldots,K\}$. For $i=0$, we have $||Q_{s,a,c_{0}}^{\pi_{t},p_{t}}||_{\infty}\leq\dfrac{1}{(1-\gamma)\lambda}\leq\min\{\xi/2,1/2\}<\dfrac{1}{1-\gamma}$ as $\xi\leq\dfrac{1}{1-\gamma}$. Hence, from (B),
@@ Appendix D Robust policy evaluator based on KL divergence
$\text{where }V_{c_{n}}^{t}(s^{\prime}):=\sum_{a^{\prime}\in\mathcal{A}}\pi(s^{\prime},a^{\prime})Q_{c_{n}}^{t}(s^{\prime},a^{\prime}).$
@@
2:$Q,V$ = Robust_Q-table($c_{i}$,$\pi$,$p^{0}$,$C_{KL}$) (see Algorithm 3)
@@
4:$T[s,s^{\prime}]=\underset{a\in\mathcal{A}}{\Sigma}\pi(a|s)P^{*}(s,a,s^{\prime}),\quad\forall(s,s^{\prime})\in\mathcal{S}\times\mathcal{S}$
5:$Q^{\pi}_{c_{i},P^{*}}=(I-\gamma T)^{-1}c_{i}$
6:$\hat{J}=\rho^{T}\left(\underset{a\in\mathcal{A}}{\Sigma}(\pi(a|s)Q^{\pi}_{c_{i},P^{*}}(s,a))\right)$
7:$d^{\pi}_{P^{*}}=(1-\gamma)(I-\gamma T)^{-1}\rho$
8:$\nabla\hat{J}=Hd^{\pi}_{P^{*}}(s)Q^{\pi}_{c_{i},P^{*}}(s,a)~{}~{}\forall~{}(s,a)\in\mathcal{S}\times\mathcal{A}$
@@
$\text{and }\theta^{*}_{(s,a)}:=\arg\min_{\theta\geq 0}\theta C_{KL}+\theta\ln{\left(\langle p^{0}(.|s,a),\exp\left(\frac{V_{c_{n}}^{t}(.)}{\theta}\right)\rangle\right)}.$
@@ E.1.2 Discussions of the result
Furthermore, Table 11 presents a comparison of wall-clock time across the algorithms. RNPG completes in the shortest time, running approximately 1.6×1.6$\times$1.6 × faster than RPPG and at least 4×4$\times$4 × faster than EPIRC_PGS (at $\gamma=0.9$). These results demonstrate that RNPG not only achieves competitive performance but also does so with significantly improved computational efficiency compared to both RPPG and EPIRC_PGS.
@@ E.2.3 Discussion of Results
In addition to performance, we compare the computational efficiency of the algorithms. Table 11 shows that RNPG requires a computation time comparable to RPPG, but significantly outperforms EPIRC_PGS in terms of speed. Specifically, RNPG is at least 5×5$\times$5 × faster than EPIRC_PGS when $\gamma=0.9$, and nearly 8×8$\times$8 × faster when $\gamma=0.995$. The increased runtime for EPIRC_PGS at higher discount factors is attributed to the longer binary search required for constraint satisfaction as $\gamma$ approaches 1. The key takeaway from this experiment is that RNPG demonstrates greater sensitivity to constraint boundaries and exhibits strong potential for scalability to larger state and action spaces. Notably, the Garnet environment used in this study contains 15 states and 20 actions. These results suggest that, with efficient implementation, RNPG can be effectively extended to high-dimensional settings.
@@ E.3.1 Environment description
$(x=0,y=0)$ up $(x=0,y=0):2/3$,$(x+1,y)=1/6$ and $(x,y+1)=1/6$
$(x=0,y\neq 0)$ up $(x=0,y):1/2$,$(x+1,y)=1/6$,$(x,y+1)=1/6$ and $(x,y-1)=1/6$
$(x\neq 0,y\neq 0)$ up $(x-1,y):1/2$,$(x+1,y)=1/6$,$(x,y+1)=1/6$ and $(x,y-1)=1/6$
$(x=0,y=0)$ left $(x=0,y=0):2/3$,$(x+1,y)=1/6$ and $(x,y+1)=1/6$
$(x\neq 0,y=0)$ left $(x,y=0):1/2$,$(x+1,y)=1/6$, $(x-1,y)=1/6$ and $(x,y+1)=1/6$
$(x\neq 0,y\neq 0)$ left $(x,y-1):1/2$,$(x+1,y)=1/6$,$(x,y+1)=1/6$ and $(x-1,y)=1/6$
$(x=3,y=3)$ down $(x=3,y=3):2/3$,$(x-1,y)=1/6$ and $(x,y-1)=1/6$
$(x=3,y\neq 3)$ down $(x=3,y):1/2$,$(x-1,y)=1/6$,$(x,y+1)=1/6$ and $(x,y-1)=1/6$
$(x\neq 3,y\neq 3)$ down $(x+1,y):1/2$,$(x-1,y)=1/6$,$(x,y+1)=1/6$ and $(x,y-1)=1/6$
$(x=3,y=3)$ right $(x=3,y=3):2/3$,$(x-1,y)=1/6$ and $(x,y-1)=1/6$
$(x\neq 3,y=3)$ right $(x,y=3):1/2$,$(x+1,y)=1/6$, $(x-1,y)=1/6$ and $(x,y-1)=1/6$
$(x\neq 3,y\neq 3)$ right $(x,y-1):1/2$,$(x+1,y)=1/6$,$(x,y+1)=1/6$ and $(x-1,y)=1/6$
@@
The rewards and cost functions are given in Table 7. In particular, if it reaches the goal state the reward is $+1$. If the agent hits the obstacle, the cost is $1$, and and the frozen grid it is $0.3$. Note that the a grid is obstacle or not is decided randomly at the beginning of an episode.
@@ E.3.2 Discussion of results
The results obtained of the given environment is depicted in Figure 6. As seen in Figure 6(b), all the three algorithms successfully learns the feasible policies. However, on observing Figure 6(a), we can clearly notice the dominance of RNPG by learning policies with better rewards. From Table 11, we see that for the frozen lake environment, the computation time of RPPG and RNPG is almost comparable. However, RNPG is atleast $3$x as faster as EPIRC_PGS for $\gamma=0.9$ and almost $4$x faster than EPIRC_PGS for $\gamma=0.995$.
@@ E.4.2 Discussion of results
In this subsection, we will present the performance of the RPPG, EPIRC_PGS and our algorithm (RNPG)(Figure 7). As shown in Figure 7(b), due to the randomness of the environment, the algorithms have some minor fluctuations. However, in this environment, RPPG and RNPG obey the constraints for the complete duration, but, EPIRC_PGS violates the constraint. Although none of the algorithms stabilize completely yet RPPG and RNPG are in the safe zone. In terms of objective return, it can be seen from Figure 7(a), the expected return for RNPG is predominantly higher than RPPG and EPIRC_PGS. While comparing the time (Table 11) for completion RPPG is the fastest in this environment marginally beating RNPG but still the speeds of both algorithms are comparable. When compared with EPIRC_PGS, RNPG is winning fairly with a speedup of $2$x compared to EPIRC_PGS when $\gamma=0.9$ and a speedup of nearly $3$x compared to EPIRC_PGS when $\gamma=0.995$.
@@ F.2 Robust Projected Policy Gradient (RPPG)
Appendix G Connection with the CRPO

### v2

@@ 1 Introduction
$\textbf{RCMDP objective:}~\underset{\pi}{\min}\underset{P\in\mathbb{P}}{\max}~J^{\pi,P}_{c_{0}}~~~\text{s.t.}~~\underset{P\in\mathbb{P}}{\max}~J^{\pi,P}_{c_{i}}\leq b,\quad i\in\{1,\ldots,K\}.$ (1)
@@ 2 Problem Formulation
$V^{\pi,P}_{c_{i}}(s)=\mathbb{E}_{P,\pi}\left[\underset{t=1}{\overset{\infty}{\Sigma}}\gamma^{t-1}\pi(a|s)c_{i}^{t}(s,a)\right],$ (4)
@@
$\text{{CMDP objective:}}~~\min~J_{c_{0}}^{\pi,P}~~\text{s.t.}~~J_{c_{i}}^{\pi,P}\leq b_{i}\quad\forall i\in\{1,\ldots,K\}.$ (5)
@@
$J^{*}=\underset{\pi\in\Pi}{\min}\underset{\lambda\in\mathbb{R}^{N}_{+}}{\max}\underset{P\in\mathbb{P}}{\max}J^{\pi,P}_{c_{0}}+\underset{i=1}{\overset{N}{\Sigma}}\lambda_{i}.\underset{P\in\mathbb{P}}{\max}(J_{c_{i}}^{\pi,P}-b_{i}).$ (8)
@@ 3.2 Policy Optimization Algorithm
$\pi_{t+1}\in\arg\min_{\pi\in\Pi}\langle\nabla_{\pi_{t}}J_{i}(\pi_{t}),\pi-\pi_{t}\rangle+\dfrac{1}{\alpha_{t}}\mathrm{KL}(\pi||\pi_{t})$ $\text{where }i=\arg\max\{\frac{J_{c_{0}}^{\pi}}{\lambda},\{J_{c_{n}}^{\pi}-b_{n}+\xi\}_{n=1}^{K}\}$
where $\mathrm{KL}$ is the usual Kullback-Leibler divergence. Note that this is a convex optimization problem, and can be optimized efficiently. If we use, $\ell_{2}$ regularization, i.e., $\pi-\pi_{t}||_{2}^{2}$, then it becomes a robust projected policy gradient (RPPG) adapted from the unconstrained version (see Appendix D[46, 36], a variant of which is used in [8] to find optimal policy for each $b_{0}$. Of course, our approach also works for $\ell_{2}$ norm which we define in Algorithm 4. Empirically, we observe that KL-divergence has a better performance, and provide iteration complexity for RNPG.
@@ 4 Theoretical Results
Thus, Algorithm 1 has an iteration complexity of $O(\epsilon^{-2})$. Also, the policy is feasible and only at most $\epsilon$-sub optimal. Our result improves upon the result $O(\epsilon^{-4})$ achieved in [8]. Further, they did not guarantee feasibility of the policy (rather, only $\epsilon$-violation). More importantly, we do not need to employ a binary search algorithm. Thus, our algorithm is computationally more efficient. Our dependence on $S$, AA, and $(1-\gamma)^{-1}$ are significantly better compared to [8] as well. Note that for the unconstrained case the iteration complexity is $O(\epsilon^{-1})$ [45]; whether we can achieve such a result for robust CMDP has been left for the future.
@@ 5 Experimental Results
We evaluate our algorithms [^2] The complete implementation is available at https://github.com/Sourav1429/RCAC_NPG on two environments: (i) Garnet, and (ii) Constrained Riverswim (CRS). Additional experimental results are provided in Appendix E.
@@ 6 Discussions and Limitation
Note that since we are not assuming strict feasibility of the optimal policy, we can only bound the violation up to $\epsilon$. The key here is to use $\lambda=2H/\epsilon$ as we do not know $\xi$, and then obtain an $\epsilon^{2}$-close result using Lemma 4.2. This makes the iteration complexity of $O(\epsilon^{-4})$. Note that our dependence on $S$, AA, and $1/(1-\gamma)$ are significantly better compared to EPIRC-PGS[8]. Further, we do not employ binary search. The proof is in Appendix C.
@@ Appendix A Proof of Proposition 1
[…]
@@
Here, we use the fact that $Q_{s,a,c_{i}}^{\pi_{t},p_{t}}||_{\infty}\leq\dfrac{1}{1-\gamma}$. This is easy to discern for $i=\{1,\ldots,K\}$. For $i=0$, we have $Q_{s,a,c_{0}}^{\pi_{t},p_{t}}||_{\infty}\leq\dfrac{1}{(1-\gamma)\lambda}\leq\min\{\xi/2,1/2\}<\dfrac{1}{1-\gamma}$ as $\xi\leq\dfrac{1}{1-\gamma}$. Hence, from (B),
@@ Appendix D Robust policy evaluator based on KL divergence
$\text{where }V_{c_{n}}^{t}(s^{\prime})=\sum_{a^{\prime}\in\mathcal{A}}\pi(s^{\prime},a^{\prime})Q_{c_{n}}^{t}(s^{\prime},a^{\prime}).$
@@
2:$Q,V$ = Robust_Q-table($c_{i}$$\pi$$p^{0}$$C_{KL}$) (see Algorithm 3)
@@
[…]
@@
$\text{and }\theta^{*}_{(s,a)}=\arg\min_{\theta\geq 0}\theta C_{KL}+\theta\ln{\left(\langle p^{0}(.|s,a),\exp\left(\frac{V_{c_{n}}^{t}(.)}{\theta}\right)\rangle\right)}.$
@@ E.1.2 Discussions of the result
Furthermore, Table 11 presents a comparison of wall-clock time across the algorithms. RNPG completes in the shortest time, running approximately $1.6\times$ faster than RPPG and at least $4\times$ faster than EPIRC_PGS (at $\gamma=0.9$). These results demonstrate that RNPG not only achieves competitive performance but also does so with significantly improved computational efficiency compared to both RPPG and EPIRC_PGS.
@@ E.2.3 Discussion of Results
In addition to performance, we compare the computational efficiency of the algorithms. Table 11 shows that RNPG requires a computation time comparable to RPPG, but significantly outperforms EPIRC_PGS in terms of speed. Specifically, RNPG is at least $5\times$ faster than EPIRC_PGS when $\gamma=0.9$, and nearly $8\times$ faster when $\gamma=0.995$. The increased runtime for EPIRC_PGS at higher discount factors is attributed to the longer binary search required for constraint satisfaction as $\gamma$ approaches 1. The key takeaway from this experiment is that RNPG demonstrates greater sensitivity to constraint boundaries and exhibits strong potential for scalability to larger state and action spaces. Notably, the Garnet environment used in this study contains 15 states and 20 actions. These results suggest that, with efficient implementation, RNPG can be effectively extended to high-dimensional settings.
@@ E.3.1 Environment description
$(x=0,y=0)$ up (x=0,y=0):2/3(x=0,y=0):2/3,$(x+1,y)=1/6$ and $(x,y+1)=1/6$
$(x=0,y\neq 0)$ up (x=0,y):1/2(x=0,y):1/2,(x+1,y)=1/6(x+1,y)=1/6,$(x,y+1)=1/6$ and (x,y1)=1/6(x,y-1)=1/6
$(x\neq 0,y\neq 0)$ up (x1,y):1/2(x-1,y):1/2,(x+1,y)=1/6(x+1,y)=1/6,$(x,y+1)=1/6$ and (x,y1)=1/6(x,y-1)=1/6
$(x=0,y=0)$ left (x=0,y=0):2/3(x=0,y=0):2/3,$(x+1,y)=1/6$ and $(x,y+1)=1/6$
$(x\neq 0,y=0)$ left (x,y=0):1/2(x,y=0):1/2,$(x+1,y)=1/6$, (x1,y)=1/6(x-1,y)=1/6 and $(x,y+1)=1/6$
$(x\neq 0,y\neq 0)$ left (x,y1):1/2(x,y-1):1/2,(x+1,y)=1/6(x+1,y)=1/6,$(x,y+1)=1/6$ and (x1,y)=1/6(x-1,y)=1/6
$(x=3,y=3)$ down (x=3,y=3):2/3(x=3,y=3):2/3,(x1,y)=1/6(x-1,y)=1/6 and (x,y1)=1/6(x,y-1)=1/6
$(x=3,y\neq 3)$ down (x=3,y):1/2(x=3,y):1/2,(x1,y)=1/6(x-1,y)=1/6,$(x,y+1)=1/6$ and (x,y1)=1/6(x,y-1)=1/6
$(x\neq 3,y\neq 3)$ down (x+1,y):1/2(x+1,y):1/2,(x1,y)=1/6(x-1,y)=1/6,$(x,y+1)=1/6$ and (x,y1)=1/6(x,y-1)=1/6
$(x=3,y=3)$ right (x=3,y=3):2/3(x=3,y=3):2/3,(x1,y)=1/6(x-1,y)=1/6 and (x,y1)=1/6(x,y-1)=1/6
$(x\neq 3,y=3)$ right (x,y=3):1/2(x,y=3):1/2,$(x+1,y)=1/6$, (x1,y)=1/6(x-1,y)=1/6 and (x,y1)=1/6(x,y-1)=1/6
$(x\neq 3,y\neq 3)$ right (x,y1):1/2(x,y-1):1/2,(x+1,y)=1/6(x+1,y)=1/6,$(x,y+1)=1/6$ and (x1,y)=1/6(x-1,y)=1/6
@@
The rewards and cost functions are given in Table 7. In particular, if it reaches the goal state the reward is +1+1. If the agent hits the obstacle, the cost is $1$, and and the frozen grid it is $0.3$. Note that the a grid is obstacle or not is decided randomly at the beginning of an episode.
@@ E.3.2 Discussion of results
The results obtained of the given environment is depicted in Figure 6. As seen in Figure 6(b), all the three algorithms successfully learns the feasible policies. However, on observing Figure 6(a), we can clearly notice the dominance of RNPG by learning policies with better rewards. From Table 11, we see that for the frozen lake environment, the computation time of RPPG and RNPG is almost comparable. However, RNPG is atleast 33x as faster as EPIRC_PGS for $\gamma=0.9$ and almost 44x faster than EPIRC_PGS for $\gamma=0.995$.
@@ E.4.2 Discussion of results
In this subsection, we will present the performance of the RPPG, EPIRC_PGS and our algorithm (RNPG)(Figure 7). As shown in Figure 7(b), due to the randomness of the environment, the algorithms have some minor fluctuations. However, in this environment, RPPG and RNPG obey the constraints for the complete duration, but, EPIRC_PGS violates the constraint. Although none of the algorithms stabilize completely yet RPPG and RNPG are in the safe zone. In terms of objective return, it can be seen from Figure 7(a), the expected return for RNPG is predominantly higher than RPPG and EPIRC_PGS. While comparing the time (Table 11) for completion RPPG is the fastest in this environment marginally beating RNPG but still the speeds of both algorithms are comparable. When compared with EPIRC_PGS, RNPG is winning fairly with a speedup of 22x compared to EPIRC_PGS when $\gamma=0.9$ and a speedup of nearly 33x compared to EPIRC_PGS when $\gamma=0.995$.
@@ F.2 Robust Projected Policy Gradient (RPPG)
Appendix G escalating to higher space
Algorithm 5 Robust Constrained Actor-Critic (RCAC)
1:Input: Environment $\mathcal{E}$, learning rates $\alpha_{V},\alpha_{C},\alpha_{\pi}$, discount factor $\gamma$, perturbation level $\delta$, constraint bound $b$, trade-off parameter $\lambda$, total episodes NN
2:Initialize actor $\pi_{\theta}$, critic $V_{\psi}$ (for reward), critic $C_{\phi}$ (for constraint)
3:for each training epoch do
4:  for each batch of size $B$ do
5:   Initialize replay buffer $\mathcal{D}\leftarrow\emptyset$
6:   Reset environment and observe initial state $s$
7:   for $i=1$ to $B$ do
8:     Sample action $a\sim\pi_{\theta}(\cdot|s)$
9:     Execute $a$ in $\mathcal{E}$; observe $r$, $c$, $s^{\prime}$, done flag
10:     Store transition $(s,a,\log\pi(a|s),r,c,s^{\prime},done)$ in $\mathcal{D}$
11:     if donedone then reset $s\leftarrow\mathcal{E}.\text{reset}()$
[…]
13:     end if
14:   end for
15:   Convert replay buffer $\mathcal{D}$ to tensors
16:   Compute advantage estimates:
$\text{adv}_{r}=r+\gamma(1-\text{done})V_{\psi}(s^{\prime})-V_{\psi}(s)$
$\text{adv}_{c}=c+\gamma(1-\text{done})C_{\phi}(s^{\prime})-C_{\phi}(s)$
17:   Compute perturbed advantages:
[…]
18:   Update critic parameters:
$\psi\leftarrow\psi-\alpha_{V}\nabla_{\psi}\mathcal{L}_{V}\text{ and }\quad\phi\leftarrow\phi-\alpha_{C}\nabla_{\phi}\mathcal{L}_{C}$
where $\mathcal{L}_{V}=\text{MSE}(r+\gamma V(s^{\prime}),V(s))$, $\mathcal{L}_{C}=\text{MSE}(c+\gamma C(s^{\prime}),C(s))$
19:   Select constraint-violating component:
$\text{choice}=\arg\max\left(\frac{V(s)}{\lambda},C(s)-b\right)$
20:   Compute actor loss:
$\mathcal{L}_{\pi}=-\log\pi(a|s)\cdot\delta_{r}\cdot(1-choice)-\log\pi(a|s)\cdot\delta_{c}\cdot(choice),$
21:   Update actor parameters:
$\theta\leftarrow\theta-\alpha_{\pi}\nabla_{\theta}\mathcal{L}_{\pi}$
22:  end for
23:end for
Appendix H Connection with the CRPO